OpenStudy (anonymous):

A company discovers that it costs $18 to produce 2 calculators, $66 to produce 4 calculators, and $450 to produce 10 calculators. Using quadratic functions, find the total cost of producing 8 calculators.

OpenStudy (anonymous):

yes. i just checked over it again to make sure.

OpenStudy (amistre64):

(2,18) (4,66) (10,450) b = (18-66)(100-4) + (450-18)(16-4) ------------------------------ (2-4)(100-4) + (10-2)(16-4)

OpenStudy (amistre64):

b = (-48)(96) + (432)(12) ------------------- (-2)(96) + (8)(12)

OpenStudy (amistre64):

b = -4608 + 5184 576 ------------- = ---- = -6 -192 + 96 -96 b = -6

OpenStudy (amistre64):

(2,18) (4,66) (10,450) a = (66-18) (4-2)(6) ------- - ----- (16-4) (16-4) a = 48 12 36 ---- - ----- = ----- = 3 12 12 12 a = 3; b = -6

OpenStudy (amistre64):

c = 18 - 4(3) - 2(-6) c = 18 -12 +12 = 18 3x^2 -6x +18 let me double check that

OpenStudy (anonymous):

i really have no idea how you remember all those steps!! this would so much easier with a graphing calc!

OpenStudy (amistre64):

lol... thats ok, i have to redo it :) im gonna work it out on the paper, itll just be easier for me :)

OpenStudy (anonymous):

Graphing calculators are for babies.

OpenStudy (anonymous):

graphing calculators are for people who really don't like math & like to take the easy way to get it done with faster.

OpenStudy (anonymous):

well if i would've found that or someone would've gave me the site, i would use it. but they didn't so i couldn't.

OpenStudy (anonymous):

amistre... ready?

OpenStudy (amistre64):

its 1:45 am here and my brain aint behaving itself ....sorry racer, maybe tomorrow :)

OpenStudy (anonymous):

this is the last question & i'll be done..

OpenStudy (anonymous):

Do you want me to do it?

OpenStudy (anonymous):

well it's too late now. i thought i would be able to figure out the answer from what she gave me. but it was wrong & i could only miss 1 out of 5.

OpenStudy (anonymous):

OK... later

OpenStudy (anonymous):

thanks though. guess i'll try again tomorrow.

OpenStudy (amistre64):

After some sleep, I see where I went wrong :) b=-6 is good, but when I used it down here I made it "+6".....which of course threw off my numbers.... (2,18) (4,66) (10,450) a = (66-18) (4-2)(-6) <-- right there ------- - ----- (16-4) (16-4) a = 48 -12 60 ---- - ----- = ----- = 5 12 12 12 a = 5; b = -6 ; now THOSE are the right numers :)

OpenStudy (amistre64):

c = 18 -4(a) - 2(b) c = 18 - 4(5) - 2(-6) c = 18 -20 +12 c = 30 - 20 c = 10 y = 5x^2 -6x +10 is the good equations. y = 5(8)^2 -6(8) +10 y = 5(64) -48 +10 y = 320 -38 y = 282 the cost of 8 calculator costs $282